### v2

@@ Abstract
Individual behavioural heterogeneity is key to understanding how natural populations adapt flexibly to environmental changes. In ant species like Aphaenogaster senilis, two key roles in collective foraging are commonly recognized: scouts, who discover food patches, and recruits, who exploit these patches and transport food back to the nest. These roles involve distinct movement patterns and exploratory behaviours. In this chapter, we develop a correlated random walk model on a bounded honeycomb lattice to interpret and replicate empirical observations of foraging ants in an enclosed arena with honeycomb tiling. We do so by extending the theory of first-passage processes for $\mathcal{N}$ random walkers when individuals belong to a heterogeneous population. We apply this theory to examine how individual behavioural heterogeneity in ants affects collective foraging efficiency, focusing on first-passage time statistics for nest-to-patch and patch-to-patch movements. With the combined use of the mathematical model and the controlled experimental setup we evaluate (i) the impact of distinct movement strategies by scouts and recruits on finding food items and (ii) whether ants practice strict central place foraging or utilise previously discovered patches as starting points for further exploration.
@@ 2 Experimental protocol in the Y-maze arena
Foraging experiments were carried out with Aphaenogaster senilis, a Mediterranean ant species displaying flexible foraging strategies, ranging from individual to group recruitment Cerda2009 . To reproduce the natural spatiotemporal scales at which these ants forage, we recorded three hours of video per trial, capturing their movements within a $2\times 1\,$ meters squared honeycomb-like arena (see Figure 1 in ref. cristin2024spatiotemporal ). This Y-maze pattern is widely utilized in studies of animal decision-making (e.g., Czaczkes2018 ), and simplifies the quantification of animal behaviour and movement since at each vertex the walker has only three choices: to go left, right or backwards. Additionally, spatial discreteness is well-suited for incorporating the present random walk theory, network theory and other statistical physics frameworks such as spin glasses Cristin2020 ; cristin2024spatiotemporal .
@@
We tracked individual ant trajectories from the videos at a temporal resolution of 2 Hz. We analysed several aspects of their movement and classified them into scouts or recruits according to the tracks’ average distance to the nest. After, the tracks were discretised and each position was assigned to the closest node (Y-shaped intersection) in the lattice. We then estimated the turning probabilities at 1-second intervals, computing whether the ant remained in the same node ($c$), or turned to the right ($r$), left ($l$), or backward ($b$), relative to its last position. This was performed separately for scouts and recruits. The relevant parameters for the recruits are $r\approx 0.0818$, $b\approx 0.0539$, $l\approx 0.0821$, and $c\approx 0.7822$, while for the scouts we found $r\approx 0.1039$, $b\approx 0.0358$ and $l\approx 0.1008$, and $c\approx 0.7595$, where all movement parameters are rounded to 4 decimal places. In Figure 1 we represent some example tracks for both scouts and recruits, to showcase their different movement patterns.
@@ 4.2 Mean Number of Distinct Visited Sites
where we have used the property that the sum of the occupation probability over all the sites and states equals unity to recover the well-known form of the MNDVS in the second line blumen1981energy ; rws_on_latticesII . For non-isotropic lattices, e.g., in the presence of reflecting domains, one needs to employ Eq. (12) contrary to what is erroneously implied in ref. LucaPRX .
@@ 5 Search of $\mathcal{N}$ distinct walkers
We first present general theoretical findings in Sec. 5.1 and 5.1 pertaining to search with $\mathcal{N}$ independent random walkers, while in Sec. 5.2 we consider model predictions pertaining to the ant experiment. Specifically, we consider the fFPT, which is the probability that the first walker in the population reaches any target of a set $\mathcal{T}$ for the first time at time $t$, in a heterogeneous population with individuals displaying different movement characteristics. The exact theoretical development for the fFPT that we present below is particularly relevant now given the recent surge in closed-form expressions for the first-passage and splitting probability generating function in a range of Markovian LucaPRX ; giuggioli2022spatio ; sarvaharmangiuggioli2023 and non-Markovian marris2024persistent walks. From the biological perspective, while in general ants share information in the food gathering process AntEcology_chapter_foraging , the application of the fFPT formalism is pertinent to our experiments as recent work on A. senilis has elucidated that during the exploration phase, i.e., before any food is found, individual ants move largely independent of each other cristin2024spatiotemporal , justifying the assumption of independent walkers.
@@ 5.1 Theory of first first-passage processes in a heterogeneous population with multiple targets
Most studies of the fFPT with $\mathcal{N}$ walkers have focused on large homogeneous populations of walkers searching in unbounded space lawley2020universal ; lawley2020probabilistic ; meerson2015mortality ; weiss1983order . Here instead, guided by the observations of A. senilis, we consider a small population of $\mathcal{N}$ ($\mathcal{N}\leq 10$) individuals with unequal movement statistics. As an illustrative example we consider a search process in bounded periodic space and study two cases: a population comprised of both strongly persistent and diffusive walkers on the square lattice, and a population of correlated walkers, representing scouts and recruits, on a honeycomb lattice.
@@
$\widetilde{\bm{P}}_{\bm{n}_{0},m_{0}}^{(p)}(\bm{n},z)=\frac{1}{N^{2}}\sum_{k_{1}=0}^{N-1}\sum_{k_{2}=0}^{N-1}\text{e}^{-\frac{2\pi i}{N}\bm{k}^{\intercal}\cdot[\bm{n}-\bm{n}_{0}]}\left[\mathbb{I}-z\bm{\lambda}\left(\frac{2\pi k_{1}}{N},\frac{2\pi k_{2}}{N}\right)\right]^{-1}\cdot U_{m_{0}},$ (19)
@@
where $\mathbb{H}(z)_{l,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{l},m_{\bm{s}_{l}},z)$, $\mathbb{H}(z)_{k,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{k},m_{\bm{s}_{k}},z)$ and $\mathbb{H}^{(i)}(\bm{n}_{0},m_{0},z)$ is the same, but with the $i^{\text{th}}$ column replaced with $\mathdutchcal{F}(n_{0},m_{0_{j}},z)=\alpha_{m_{0_{j}}}\left[\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}(\bm{s}_{1},m_{\bm{s}_{1}},z),...,\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}(\bm{s}{N},m_{\bm{s}_{\mathdutchcal{N}}},z)\right]^{\intercal}$. The splitting probability for the Markovian diffusive case may be found in Eq. (11) of ref. giuggioli2022spatio . The notation $\alpha_{m_{0_{j}}}P_{\bm{n}_{0},m_{0_{j}}}(\bm{n},m,t)=\alpha_{m_{0_{j}}}\bm{e}_{m}^{\intercal}\cdot\bm{P}_{\bm{n}_{0}}(\bm{n},t)\cdot\bm{e}_{m_{0_{j}}}$, is the occupation probability that the walker occupies site $\bm{n}$ and state $m$ given the initial condition $(\bm{n}_{0},m_{0_{j}})$, which is weighted by $\alpha_{m_{0_{j}}}$. Presently, due to the assumption of each direction being equi-probable at time zero $\alpha_{i}=1/3$, $\forall i$. Moreover, since each site can be accessed via three distinct internal states, to represent a fully absorbing trap we set the probability of leaving each internal state in the trap site to zero. For example, to place a target in a positive site, absorbing traps in the states $1,2$ and $3$ are required. Hence, for all cases presented here, and elsewhere in this chapter, the target set $\mathcal{T}$ is comprised of sets of three target states for each physical site. Henceforth, for ease of notation, when discussing specific cases we denote the set of targets with only the physical location(s) at site(s) $\bm{s}_{i}$.
@@
Figure 4: The fFPT for different heterogeneous populations of random walkers on the periodic square lattice. The population is comprised of $\mathcal{N}=\mathcal{N}_{\text{corr}}+\mathcal{N}_{\text{diff}}=8$ random walkers. In panel (a) we evaluate Eq. (18) with the first-passage terms replaced with the splitting probability for the single persistent walker given by Eq. (21). The population starts at $\bm{n}_{0}=(10,10)$ and is searching for a nearby target at $\bm{s}=(7,7)$ and a far away target at $\bm{s}=(1,9)$ in a domain size $N_{1}\times N_{2}=20\times 20$. Hence, with the square lattice requiring four internal states to represent the possible previous movement directions, and the system contains two physical target sites, we have $\mathdutchcal{N}=8$ target states. For the diffusive population we take $q=0.95$, while the persistent population has a persistence length $p_{l}=20$. The inset shows the overall probability that each target is found before the other (the summation of the splitting probability over time) as a function of $\mathcal{N}_{\text{corr}}$ for the same parameter regime. Panel (b) shows the MfFPT, $\mathcal{F}_{\bm{n}_{0}}^{(\mathfrak{h})}(\mathcal{T};\{1,\mathcal{N}\}=\sum_{t}F_{\bm{n}_{0}}^{(\mathfrak{h})}(\mathcal{T},t;\{1,\mathcal{N}\})$, as a function of $\mathcal{N}_{\text{corr}}$ with varying persistent lengths all chosen past the optimum for this system $p_{l}\approx 4.76$. Note that the diffusive persistence length is $p_{l}=4(3q)^{-1}$ and for the diffusive members of the population we take $q=0.95$ ($p_{l}\approx 1.40)$. In this case we choose a domain size of $N_{1}\times N_{2}=14\times 14$, $\bm{n}_{0}=(7,7)$ and $\bm{s}=(14,14)$ meaning $\mathdutchcal{N}=4$. Dots represent the result of $10^{6}$ stochastic simulations.
@@ 5.2 Experimentally relevant search on the Y-maze arena
To implement a reflecting boundary condition on the honeycomb lattice, note that although there are positive and negative states occupying the unit cells that lie at the edge of the domain, if the site is not at the corner, only one of the two states will interact with the boundary creating the so-called zig-zag boundary condition henry2003random ; marris2023exact . To elucidate, consider Fig. 2 (iii) where, although the positive site labelled y, lies in unit cell at the edge of the square lattice, all three movement options are available to the walker leaving the site. In contrast, travelling horizontally East from the negative site labelled e would cause a reflection to occur. In general, for the $\bm{n}=(N_{1},n_{2})$ boundary, the walker may only hit the boundary if it leaves a negative site, while at the $\bm{n}=(n_{1},N_{2})$ boundary it is the positive site that interacts with the boundary, meaning at $\bm{n}=(N_{1},N_{2})$, both sites are part of the reflecting boundary.
To implement this zig-zag reflecting boundary condition via a fully analytical procedure is a formidable task when the movement is not diffusive. To proceed we use a semi-analytical iterative procedure gueneau2024siegmund ; marris2024persistent , which has been used to study persistent motion with complicated boundary conditions where the analytical representation of the occupation probability is not available. This procedure requires the decomposition of the movement dynamics into matrices that govern only the jumps from each state $i$ to every other state $j$, denoted by $\mathbb{M}(i,j)$. In our case this leads to 36 sparse transition matrices, of size $N_{1}N_{2}\times N_{1}N_{2}$. More concretely, the matrices that govern movement in the bulk have $1/N_{1}N_{2}$ ($\approx 0.31\%$) non-zero elements, while those governing movement on the boundary have $1/N_{1}^{2}N_{2}$ ($\approx 0.022\%$) or $1/N_{1}N_{2}^{2}$ ($\approx 0.0135\%$) non-zero elements. Hence, all matrices benefit from the use of sparse methods leading to efficient computation. See Appendix 3 for a detailed discussion of the transition matrices.
@@ 6 Results
With ant colonies facing scenarios with multiple food sources in nature, we have explored how switching behaviour, e.g., from scout to recruit, may affect search efficiency in a landscape with multiple targets (Fig. 7). For each experiment we assume a homogeneous population comprised of the number $\mathcal{N}$ ants observed experimentally in the laboratory set-up (written explicitly in Fig. 6), starting either as scouts or recruits. By assuming that the entire population may modify searching patterns upon finding the first target or when returning back to the nest, we address the question of behavioural switching in two possible foraging strategies; (i) nest-to-patch foraging, i.e., strict central place foraging, in which ants go back to the nest before searching for the next target, and (ii) patch-to-patch foraging, where the ants use the previously discovered patch as a starting point to search for new food targets. The theoretical time for scenario (i) is defined as $\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{2}};\{1,\mathcal{N}\})=\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{1}};\{1,\mathcal{N}\})+\mathfrak{F}_{\bm{s_{1}}}(\bm{n}_{0};\{1,\mathcal{N}\})+\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{2}};\{1,\mathcal{N}\})$, with $\bm{s}_{1}$ being the first found patch, and $\bm{s}_{2}$ the second one, while we define the theoretical time for scenario (ii) as $\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{2}};\{1,\mathcal{N}\})=\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{1}};\{1,\mathcal{N}\})+\mathfrak{F}_{\bm{s_{1}}}(\bm{s_{2}};\{1,\mathcal{N}\})$. Note that since the process is bounded by reflecting walls, the first first-passage probability is not symmetric with respect to the initial condition and target i.e., $F_{\bm{n}_{0}}(\bm{s},t;\{1,\mathcal{N}\})\neq F_{\bm{s}}(\bm{n}_{0},t;\{1,\mathcal{N}\})$.
@@
We calculated the Euclidean distance between the mode of the theoretical distribution and the corresponding first capture time in each empirical setup. The results shown in Fig. 8 indicate that A. senilis likely uses a central-place foraging strategy (i.e., Nest to Patch), with scouts initially searching for the first patch. The results are similar regardless of whether a behavioural switch from scout to recruit (Sc.-Re.) occurs after a return to the nest after finding the first patch or whether scouts continue the search to the second patch (Sc.-Sc.). Nonetheless, the trend observed in Fig. 8 is not statistically supported by pairwise Wilcoxon tests, and the figure highlights a heterogeneous distribution of data points as certain setups reveal close alignment between theoretical and experimental behaviours, while others do not. This is consistent with the analyses carried out in Figs. 6 and 7.

### v3

@@ Abstract
Behavioral heterogeneities in animals, also known as syndromes, play a crucial role in understanding how natural populations flexibly adapt to environmental changes. In ant species like Aphaenogaster senilis, two key roles in collective foraging are commonly recognised: scouts, who discover food patches, and recruits, who exploit these patches and transport food back to the nest. These roles involve distinct movement patterns and exploratory behaviours. In this chapter, we develop a correlated random walk model on a bounded honeycomb lattice to interpret and replicate empirical observations of foraging ants in an enclosed arena with honeycomb tiling. We do so by extending the theory of first-passage processes for $\mathcal{N}$ random walkers when individuals belong to a heterogeneous population. We apply this theory to examine how individual behavioural heterogeneity in ants affects collective foraging efficiency, focusing on first-passage time statistics for nest-to-patch and patch-to-patch movements. With the combined use of the mathematical model and the controlled experimental setup we evaluate (i) the impact of distinct movement strategies by scouts and recruits on finding food items and (ii) whether ants practice strict central place foraging or utilise previously discovered patches as starting points for further exploration.
@@ 2 Experimental protocol in the Y-maze arena
Foraging experiments were carried out with Aphaenogaster senilis, a Mediterranean ant species displaying flexible foraging strategies, ranging from individual to group recruitment Cerda2009 . To reproduce the natural spatiotemporal scales at which these ants forage, we recorded three hours of video per trial, capturing their movements within a $2\times 1\,$ meters squared honeycomb-like arena (see Figure 1 in ref. cristin2024spatiotemporal ). This Y-maze pattern is widely utilized in studies of animal decision-making (e.g., Czaczkes2018 ), and simplifies the quantification of animal behaviour and movement since at each vertex the walker has only three choices: to go left, right or backwards. Additionally, spatial discreteness of the arena lends itself to the use of the present random walk theory, and other statistical physics frameworks such as spin glasses cristin2024spatiotemporal ; Cristin2020 .
@@
We tracked individual ant trajectories from the videos at a temporal resolution of 2 Hz. We analysed several aspects of their movement and classified them into scouts or recruits according to the tracks’ average distance to the nest. After, the tracks were discretised in time at 1-second intervals, each position was assigned to the closest node (Y-shaped intersection) in the lattice. We then estimated the turning probabilities, computing whether the ant remained in the same node ($c$), or turned to the right ($r$), left ($l$), or backward ($b$), relative to its last position. This was performed separately for scouts and recruits. The relevant parameters, averaged over all experiments, for the recruits are $r\approx 0.0818$, $b\approx 0.0539$, $l\approx 0.0821$, and $c\approx 0.7822$, while for the scouts we found $r\approx 0.1039$, $b\approx 0.0358$ and $l\approx 0.1008$, and $c\approx 0.7595$, with all movement parameters rounded to 4 decimal places. In Figure 1 we represent some example tracks for both scouts and recruits, to showcase their different movement patterns.
@@ 4.2 Mean Number of Distinct Visited Sites
where we have used, in the second line, the property that the sum of the occupation probability over all the sites and states equals unity to recover the well-known form of the MNDVS rws_on_latticesII ; blumen1981energy . For non-isotropic lattices, e.g., in the presence of reflecting domains, one needs to employ Eq. (12) contrary to what is erroneously implied in ref. LucaPRX .
@@ 5 Search of $\mathcal{N}$ distinct walkers
We first present general theoretical findings in Sec. 5.1 and 5.1 pertaining to search with $\mathcal{N}$ independent random walkers, while in Sec. 5.2 we consider model predictions pertaining to the ant experiment. Specifically, we consider the fFPT, which is the probability that the first walker in the population reaches any target of a set $\mathcal{T}$ for the first time at time $t$, in a heterogeneous population with individuals displaying different movement characteristics. The exact theoretical development for the fFPT that we present below is particularly relevant now given the recent surge in closed-form expressions for the first-passage and splitting probability generating function in a range of Markovian LucaPRX ; sarvaharmangiuggioli2023 ; giuggioli2024multi and non-Markovian marris2024persistent walks. From the biological perspective, while in general ants share information in the food gathering process AntEcology_chapter_foraging , the application of the fFPT formalism is pertinent to our experiments as recent work on A. senilis has elucidated that during the exploration phase, i.e., before any food is found, individual ants move largely independent of each other cristin2024spatiotemporal , justifying the assumption of independent walkers.
@@ 5.1 Theory of first first-passage processes in a heterogeneous population with multiple targets
Most studies of the fFPT with $\mathcal{N}$ walkers have focused on large homogeneous populations of walkers searching in unbounded space weiss1983order ; lawley2020universal ; lawley2020probabilistic ; meerson2015mortality . Here instead, guided by the observations of A. senilis, we consider a small population of $\mathcal{N}$ ($\mathcal{N}\leq 10$) individuals with unequal movement statistics. As an illustrative example we consider a search process in bounded periodic space and study two cases: a population comprised of both strongly persistent and diffusive walkers on the square lattice, and a population of correlated walkers, representing scouts and recruits, on a honeycomb lattice.
@@
$\widetilde{\bm{P}}_{\bm{n}_{0},m_{0}}^{(p)}(\bm{n},z)=\frac{1}{N^{2}}\sum_{k_{1}=0}^{N-1}\sum_{k_{2}=0}^{N-1}\text{e}^{-\frac{2\pi i}{N}\bm{k}\cdot[\bm{n}-\bm{n}_{0}]^{\intercal}}\left[\mathbb{I}-z\bm{\lambda}\left(\frac{2\pi k_{1}}{N},\frac{2\pi k_{2}}{N}\right)\right]^{-1}\cdot U_{m_{0}},$ (19)
@@
where $\mathbb{H}(z)_{l,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{l},m_{\bm{s}_{l}},z)$, $\mathbb{H}(z)_{k,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{k},m_{\bm{s}_{k}},z)$ and $\mathbb{H}^{(i)}(\bm{n}_{0},m_{0},z)$ is the same, but with the $i^{\text{th}}$ column replaced with $\mathdutchcal{F}(n_{0},m_{0_{j}},z)=\alpha_{m_{0_{j}}}\left[\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}(\bm{s}_{1},m_{\bm{s}_{1}},z),...,\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}(\bm{s}{N},m_{\bm{s}_{\mathdutchcal{N}}},z)\right]^{\intercal}$. The notation $\alpha_{m_{0_{j}}}P_{\bm{n}_{0},m_{0_{j}}}(\bm{n},m,t)=\alpha_{m_{0_{j}}}\bm{e}_{m}^{\intercal}\cdot\bm{P}_{\bm{n}_{0}}(\bm{n},t)\cdot\bm{e}_{m_{0_{j}}}$, is the occupation probability that the walker occupies site $\bm{n}$ and state $m$ given the initial condition $(\bm{n}_{0},m_{0_{j}})$, which is weighted by $\alpha_{m_{0_{j}}}$. Presently, due to the assumption of each direction being equi-probable at time zero $\alpha_{i}=1/3$, $\forall i$. Moreover, since each site can be accessed via three distinct internal states, to represent a fully absorbing trap we set the probability of leaving each internal state in the trap site to zero. For example, to place a target in a positive site, absorbing traps in the states $1,2$ and $3$ are required. Hence, for all cases presented here, and elsewhere in this chapter, the target set $\mathcal{T}$ is comprised of sets of three target states for each physical site. Henceforth, for ease of notation, when discussing specific cases we denote the set of targets with only the physical location(s) at site(s) $\bm{s}_{i}$. The splitting probability for the Markovian case for the diffusive subset can be found in Eq. (11) of ref. giuggioli2022spatio .
@@
Figure 4: The fFPT for different heterogeneous populations of random walkers on the periodic square lattice. The population is comprised of $\mathcal{N}=\mathcal{N}_{\text{corr}}+\mathcal{N}_{\text{diff}}=8$ random walkers. In panel (a) we evaluate Eq. (18) with the first-passage terms replaced with the splitting probability for the single persistent walker given by Eq. (21). The population starts at $\bm{n}_{0}=(10,10)$ and is searching for a nearby target at $\bm{s}=(7,7)$ and a far away target at $\bm{s}=(1,9)$ in a domain size $N_{1}\times N_{2}=20\times 20$. Hence, with the square lattice requiring four internal states to represent the possible previous movement directions, and the system contains two physical target sites, we have $\mathdutchcal{N}=8$ target states. For the diffusive population we take $q=0.95$, while the persistent population has a persistence length $p_{l}=20$. The inset shows the overall probability that each target is found before the other (the summation of the splitting probability over time) as a function of $\mathcal{N}_{\text{corr}}$ for the same parameter regime. Panel (b) shows the MfFPT, $\mathcal{F}_{\bm{n}_{0}}^{(\mathfrak{h})}(\mathcal{T};\{1,\mathcal{N}\}=\sum_{t}\,tF_{\bm{n}_{0}}^{(\mathfrak{h})}(\mathcal{T},t;\{1,\mathcal{N}\})$, as a function of $\mathcal{N}_{\text{corr}}$ with varying persistent lengths all chosen past the optimum for this system $p_{l}\approx 4.76$. Note that the diffusive persistence length is $p_{l}=4(3q)^{-1}$ and for the diffusive members of the population we take $q=0.95$ ($p_{l}\approx 1.40)$. In this case we choose a domain size of $N_{1}\times N_{2}=14\times 14$, $\bm{n}_{0}=(7,7)$ and $\bm{s}=(14,14)$ meaning $\mathdutchcal{N}=4$. Dots represent the result of $10^{6}$ stochastic simulations.
@@ 5.2 Experimentally relevant search on the Y-maze arena
To implement a reflecting boundary condition on the honeycomb lattice, note that although there are positive and negative states occupying the unit cells that lie at the edge of the domain, if the site is not at the corner, only one of the two states will interact with the boundary creating the so-called zig-zag boundary condition marris2023exact ; henry2003random . To elucidate, consider Fig. 2 (iii) where, although the positive site labelled y, lies in unit cell at the edge of the square lattice, all three movement options are available to the walker leaving the site. In contrast, travelling horizontally East from the negative site labelled e would cause a reflection to occur. In general, for the $\bm{n}=(N_{1},n_{2})$ boundary, the walker may only hit the boundary if it leaves a negative site, while at the $\bm{n}=(n_{1},N_{2})$ boundary it is the positive site that interacts with the boundary, meaning at $\bm{n}=(N_{1},N_{2})$, both sites are part of the reflecting boundary.
To implement this zig-zag reflecting boundary condition via a fully analytical procedure is a formidable task when the movement is not diffusive. To proceed we use a semi-analytical iterative procedure marris2024persistent ; gueneau2024siegmund , which has been used to study persistent motion with complicated boundary conditions where the analytical representation of the occupation probability is not available. This procedure requires the decomposition of the movement dynamics into matrices that govern only the jumps from each state $i$ to every other state $j$, denoted by $\mathbb{M}(i,j)$. In our case this leads to 36 sparse transition matrices, of size $N_{1}N_{2}\times N_{1}N_{2}$. More concretely, the matrices that govern movement in the bulk have $1/N_{1}N_{2}$ ($\approx 0.31\%$) non-zero elements, while those governing movement on the boundary have $1/N_{1}^{2}N_{2}$ ($\approx 0.022\%$) or $1/N_{1}N_{2}^{2}$ ($\approx 0.0135\%$) non-zero elements. Hence, all matrices benefit from the use of sparse methods leading to efficient computation. See Appendix 3 for a detailed discussion of the transition matrices.
@@ 6 Results
With ant colonies facing scenarios with multiple food sources in nature, we have explored how switching behaviour, e.g., from scout to recruit, may affect search efficiency in a landscape with multiple targets (Fig. 7). For each experiment we assume a homogeneous population comprised of $\mathcal{N}$ ants observed experimentally in the laboratory set-up (written explicitly in Fig. 6), starting either as scouts or recruits. By assuming that the entire population may modify searching patterns upon finding the first target or when returning back to the nest, we address the question of behavioural switching in two possible foraging strategies; (i) nest-to-patch foraging, i.e., strict central place foraging, in which ants go back to the nest before searching for the next target, and (ii) patch-to-patch foraging, where the ants use the previously discovered patch as a starting point to search for new food targets. The theoretical time for scenario (i) is defined as $\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{2}};\{1,\mathcal{N}\})=\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{1}};\{1,\mathcal{N}\})+\mathfrak{F}_{\bm{s_{1}}}(\bm{n}_{0};\{1,\mathcal{N}\})+\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{2}};\{1,\mathcal{N}\})$, with $\bm{s}_{1}$ being the first found patch, and $\bm{s}_{2}$ the second one, while we define the theoretical time for scenario (ii) as $\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{2}};\{1,\mathcal{N}\})=\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{1}};\{1,\mathcal{N}\})+\mathfrak{F}_{\bm{s_{1}}}(\bm{s_{2}};\{1,\mathcal{N}\})$. Note that since the process is bounded by reflecting walls, the first first-passage probability is not symmetric with respect to the initial condition and target i.e., $F_{\bm{n}_{0}}(\bm{s},t;\{1,\mathcal{N}\})\neq F_{\bm{s}}(\bm{n}_{0},t;\{1,\mathcal{N}\})$.
@@
We calculated the Euclidean distance between the mode of the theoretical distribution and the corresponding first capture time in each empirical setup. The results shown in Fig. 8 indicate that A. senilis likely uses a central-place foraging strategy (i.e., Nest to Patch), with scouts initially searching for the first patch. The results are similar regardless of whether a behavioural switch from scout to recruit (Sc.-Re.) occurs after a return to the nest following the discovery of the first patch or whether scouts continue the search to the second patch (Sc.-Sc.). Nonetheless, the trend observed in Fig. 8 is not statistically supported by pairwise Wilcoxon tests, and the figure highlights a heterogeneous distribution of data points, that is certain setups reveal close alignment between theoretical and experimental behaviours, while others do not. This is consistent with the analyses carried out in Figs. 6 and 7.